# Privacy-Preserving Intrusion Detection in Software-defined VANET using Federated Learning with BERT

### 14 Jan 2024 v1

@@ A. Dataset: VeReMi
A benchmark dataset for misbehavior detection systems for VANETs is the VeReMi dataset. It is a synthetic dataset made with LuST (Version 2) and VEINS (with modifications based on Version 4.6). In this dataset, onboard unit message logs from a simulation environment are included, along with tagged ground truth. It has two primary uses: it gives a starting point for determining how well misbehavior detection systems work city-wide. This database has very few attack classes and sizes, as shown in Table. [13] [23]. As shown in Table, we create a balanced subset of preferred samples across all attack types for model training and evaluation.
@@ C. FL-BERT Model
The model being discussed in this paper is a practical application of federated learning, specifically utilising the BERT (Bidirectional Encoder Representations from Transformers) architecture. This model has been specifically tailored for use with the VeReMi dataset. Federated learning is an approach in machine learning where the training process is distributed over several devices or servers, referred to as clients, without the need to share the raw data. This approach guarantees the confidentiality of records, as the most efficient iterations of changes are transmitted to a central server, which consolidates these updates to enhance a global model.
Data preparation plays a key role in the pipeline of every machine learning system. In the present model, a combination of various columns extracted from the dataset is performed, including ’sendtime’, ’sender’, ’messageID’, ’pos’, ’spd’, ’sendtime’, and ’sender’, resulting in the creation of a unified column referred to as ’textual content’. This compiled text serves as the input for the BERT model. The column labeled ’AttackerType’, which potentially contains explicit data indicating unique categories of attackers is subsequently encoded using label encoding to convert it into a numerical format. The column labeled ”labels” in this encoding acts as the dependent variable for the model. The data is subsequently divided into educational and testing components, with 80% of the data allocated for educational purposes and the remaining 20% set aside for testing.

### 17 Jan 2024 v2

@@ A. Dataset: VeReMi
A benchmark dataset for misbehavior detection systems for VANETs is the VeReMi dataset. It is a synthetic dataset made with LuST (Version 2) and VEINS (with modifications based on Version 4.6). In this dataset, onboard unit message logs from a simulation environment are included, along with tagged ground truth. It has two primary uses: it gives a starting point for determining how well misbehavior detection systems work city-wide. This database has very few attack classes and sizes, as shown in Table. [12] [18]. Table shows that we create a balanced subset of preferred samples across all attack types for model training and evaluation.
@@ C. FL-BERT Model
The model being discussed in this paper is a practical application of federated learning, specifically utilising the BERT (Bidirectional Encoder Representations from Transformers) architecture. This model has been specifically tailored for use with the VeReMi dataset. Federated learning is an approach in machine learning where the training process is distributed over several devices or servers, referred to as clients, without the need to share the raw data. This approach guarantees the confidentiality of records, as the most efficient iterations of changes are transmitted to a central server, consolidating these updates to enhance a global model.
Data preparation plays a key role in the pipeline of every machine learning system. In the present model, a combination of various columns extracted from the dataset is performed, including ’sendtime’, ’sender’, ’messageID’, ’pos’, ’spd’, ’sendtime’, and ’sender’, resulting in the creation of a unified column referred to as ’textual content’. This compiled text serves as the input for the BERT model. The column labeled ’AttackerType’, potentially containing explicit data indicating unique categories of attackers, is subsequently encoded using label encoding to convert it into a numerical format. The column labelled ”labels” in this encoding acts as the dependent variable for the model. The data is subsequently divided into educational and testing components, with 80% of the data allocated for educational purposes and the remaining 20% set aside for testing.